OpenStudy (anonymous):

How would I solve this? f(x) = 1/2x + 3 and f(g) = -4x - 1 Find: g[f(4)]

hartnn (hartnn):

first try to find f(4) can you ?

hartnn (hartnn):

f(4) is when you put x=4 in f(x)

OpenStudy (anonymous):

f(4) = 1/12 put this into g(1/12)

OpenStudy (anonymous):

I know that we find f(4) first, I'm curious as to WHY we find it? Can someone give me a little background?

hartnn (hartnn):

g (f(4)) means value of g(x) AT x= f(4) for that we will need f(4) first, right ?

OpenStudy (anonymous):

u can use g(x)= f^-1(-4x-1) and f^-1(x) is 1-3x/2x replace x by -4x-1 u'll get \[g(x)= 2+6x/-4x-1\]..........(1) now f(4)= 1/11 so u can find from (1) g(1/11) which would be the answer